# Rotating black holes in de Rham-Gabadadze-Tolley massive gravity: Analytic calculation procedure

### v2

@@ Abstract
I Introduction
@@
The introduction of a mass term for gravity presents a formidable challenge. The linear mass term is plagued by the van Dam-Veltman-Zakharov (vDVZ) discontinuity [1]. On the other hand, the nonlinear mass term may propagate the sixth mode on a curved background, known as the Boulware-Deser (BD) ghost [2]. To circumvent these issues, Comelli and his cooperators try to introduce mass terms without Lorentz symmetry [3, 4]. We presented a detailed study of the spherically symmetric solutions in Lorentz breaking massive gravity [5]. Concurrently, de Rham, Gabadadze, and Tolley proposed a healthy Lorentz invariant massive gravity model [6] in 2011.
@@
In dRGT theory, the leading contribution to the decoupling limit is a total derivative which prevents the emergence of ghosts. Beyond the decoupling limit, the lapse function is a Lagrange multiplier that constraint the sixth mode in an arbitrary background. Thus, the dRGT theory described a massive spin-2 particles that propagate 5 healthy degrees of freedom. A comprehensive review of massive gravity is available in [7]. If graviton possesses mass, the gravitational waves would travel at a speed different from light. The recent observation of GW150914 by LIGO [8] has placed an upper bound on the graviton mass $m_{g}\leq 1.2\times 10^{-22}eV$. For more about the graviton mass bounds, you may see Ref. [9]. Fortunately, the dRGT theory can produce a small graviton mass technically natural [10].
@@
Numerous studies have explored black hole solutions within the dRGT theory. Nieuwenhuizen obtained the Schwarzschild-de Sitter and Reissner-Nordström-de Sitter black holes [11]. Cai et. al. also derived the static charged black hole solutions [12]. De Rham and his colleagues [13] highlighted the existence of a new basic invariant $I^{ab}=g^{\alpha\beta}\partial_{\alpha}\phi^{a}\partial_{\beta}\phi^{b}$, whose singularities could pose a problem for fluctuations. In the unitary gauge, it is evident that $I^{ab}=g^{\alpha\beta}\delta_{\alpha}^{a}\delta_{\beta}^{b}$ would exhibit a singularity if $g_{\alpha\beta}$ contains an event horizon. Consequently, there are no self-consistent static spherically symmetric (SSS) solutions in the unitary gauge. In our previous work [14], we utilized the nonunitary gauge to search for SSS solutions and discovered seven solutions, including black hole with the Stückelberg hair. We also derive the Vaidya solution in Ref. [15] and investigated the superradiant instabilities for charged furry black holes in dRGT theory [16]. Many papers have also studied the dRGT theory with degenerate reference metrics [17, 18, 19, 20, 21, 22]. And Ref. [23] calculate the SSS solutions in arbitrary dimensions.
@@
Recently, we extended Chandrasekhar’s method of calculating rotating black holes into $f(R)$ theory [24, 25, 26]. In this paper, we further extend this analytical method into the dRGT massive gravity. We proved that arbitrary nonunitary gauge $\phi^{a}$ with the Minkowski reference metric $\eta_{ab}$ equal to the unitary gauge $\bar{\phi}^{a}=x^{\alpha}\delta^{a}_{\alpha}$ with a curved reference metric $\bar{f}_{ab}$. Thus, instead of chosen the ansätz of the Stückelberg field $\phi^{a}$, we directly choose the axisymmetric reference metric $f_{ab}$ with a unitary gauge $\phi^{a}=x^{\alpha}\delta^{a}_{\alpha}$. The field equations with the electromagnetic fields can be simplified to the Ernst equations. The shortest and simplest route to obtain rotating solutions is to use the conjugate variables. We solve the homogeneous equations and the non-homogeneous equations with the conjugate variables respectively. And the solution that beyond the Kerr-Newman family is obtained.
@@
This paper is organized as follows: Sec. II review the field equation of dRGT theory [6]. In Sec. III, we give a general stationary axisymmetric metric and calculate the Einstein tensor $G_{ab}$. In Sec. IV, we discussed the electromagnetic field equations in the stationary axisymmetric background and the energy-momentum of electromagnetic fields. In Sec. V, we proved that arbitrary nonunitary gauge $\phi^{a}$ with the Minkowski reference metric $\eta_{ab}$ equal to the unitary gauge $\bar{\phi}^{a}=x^{\alpha}\delta^{a}_{\alpha}$ with a curved reference metric $\bar{f}_{ab}$. And we derive the matrix square root $\gamma$ by using the Cayley-Hamilton theorem and calculate the $T^{(\mathcal{K})}_{ab}$. In Sec. VI, we simplify the metric to a highly symmetric form by choosing the gauge freedom. In Sec. VII, we derive the Ernst equations. In Sec. VIII, we transform the variables into a conjugate metric. In Sec. IX, we solve the homogeneous equations of the Ernst equations. In Sec. X, we obtain the non-homogeneous solutions. The last section is a detailed conclusion and discussion.
@@
II The modified Einstein equations in dRGT theory
@@
And the trace of $\gamma^{i}{}^{\alpha}{}_{\beta}$ is defined as $[\gamma^{i}]=g^{\alpha\beta}\gamma^{i}_{\alpha\beta}$. In the original Ref. [6], the matrix $\gamma^{2}$ is defined by the expression
@@
and $T^{(m)}_{\alpha\beta}$ represents the energy-momentum tensor of mater. It should be pointed out that the following calculations are acually using the tetrad components $a$, not $\alpha$. Due to the Bianchi identity $\nabla_{\alpha}G^{\alpha\beta}=0$ and the conservation of energy-momentum tensor $\nabla_{\alpha}T^{(m)}{}^{\alpha\beta}=0$, there is $\nabla_{\alpha}T^{(\mathcal{K})}{}^{\alpha\beta}=0$. As is pointed in Ref.[20], the variation of $[\gamma^{i}]$ is given by the relationship
@@
$-\frac{3\alpha_{3}}{2}((6-4[\gamma]+\mathcal{U}_{2})\gamma^{\alpha}{}_{\beta}-2([\gamma]-2)\gamma^{2}{}^{\alpha}{}_{\beta}+2\gamma^{3}{}^{\alpha}{}_{\beta})$
$-2\alpha_{4}((6-6[\gamma]+3\mathcal{U}_{2}-\mathcal{U}_{3})\gamma^{\alpha}{}_{\beta}+3(2-2[\gamma]+\mathcal{U}_{2})\gamma^{2}{}^{\alpha}{}_{\beta}+6(1-[\gamma])\gamma^{3}{}^{\alpha}{}_{\beta}+6\gamma^{4}{}^{\alpha}{}_{\beta}),$ (13)
@@
III Stationary axisymmetric spacetime
@@
±$\pm e^{2\mu}[(dx^{2})^{2}+(dx^{3})^{2}],$ (28)
which is shown in Ref. [30]. At last, we shall write the stationary axisymmetric metric in the form [30]
$ds^{2}=-e^{2\nu(x^{2},x^{3})}dt^{2}+e^{2\psi(x^{2},x^{3})}(d\varphi-q(x^{2},x^{3})dt)^{2}+e^{2\mu_{2}(x^{2},x^{3})}(dx^{2})^{2}+e^{2\mu_{3}(x^{2},x^{3})}(dx^{3})^{2},$ (29)
@@
where $\mathrm{d}$ is the exterior differential operator and $\wedge$ stands for Cartan wedge. $\Omega^{a}{}_{b}$ are curvature 2-forms. After direct and tedious mathematical calculation [24, 25, 26], one obtain nonzero components of $G_{ab}$ as follows:
$G^{0}{}_{0}=$ $e^{-2\mu_{2}}\left(\psi_{,2,2}+\psi^{2}_{,2}+\psi_{,2}(\mu_{3}-\mu_{2})_{,2}\right)+e^{-2\mu_{3}}\left(\psi_{,3,3}+\psi^{2}_{,3}+\psi_{,3}(\mu_{2}-\mu_{3})_{,3}\right)$
$+e^{-\mu_{2}-\mu_{3}}\left((e^{\mu_{2}-\mu_{3}}\mu_{2,3})_{,3}+(e^{\mu_{3}-\mu_{2}}\mu_{3,2})_{,2}\right)+\frac{1}{4}e^{2(\psi-\nu-\mu_{2})}q_{,2}^{2}+\frac{1}{4}e^{2(\psi-\nu-\mu_{3})}q_{,3}^{2},$ (32)
$G^{1}{}_{1}=$ $e^{-2\mu_{2}}\left(\nu_{,2,2}+\nu^{2}_{,2}+\nu_{,2}(\mu_{3}-\mu_{2})_{,2}\right)+e^{-2\mu_{3}}\left(\nu_{,3,3}+\nu^{2}_{,3}+\nu_{,3}(\mu_{2}-\mu_{3})_{,3}\right)$
$+e^{-\mu_{2}-\mu_{3}}\left((e^{\mu_{2}-\mu_{3}}\mu_{2,3})_{,3}+(e^{\mu_{3}-\mu_{2}}\mu_{3,2})_{,2}\right)-\frac{3}{4}e^{2(\psi-\nu-\mu_{2})}q_{,2}^{2}-\frac{3}{4}e^{2(\psi-\nu-\mu_{3})}q_{,3}^{2},$ (33)
$G^{2}{}_{2}=$ $e^{-2\mu_{2}}\left(\psi_{,2}\nu_{,2}+(\psi+\nu)_{,2}\mu_{3,2}\right)+e^{-2\mu_{3}}\big{(}(\psi+\nu)_{,3,3}+\nu^{2}_{,3}+\psi^{2}_{,3}+\psi_{,3}\nu_{,3}$
$-(\psi+\nu)_{,3}\mu_{3,3}\big{)}+\frac{1}{4}e^{2(\psi-\nu-\mu_{2})}q_{,2}^{2}-\frac{1}{4}e^{2(\psi-\nu-\mu_{3})}q_{,3}^{2},$ (34)
$G^{3}{}_{3}=$ $e^{-2\mu_{3}}\left(\psi_{,3}\nu_{,3}+(\psi+\nu)_{,3}\mu_{2,3}\right)+e^{-2\mu_{2}}\big{(}(\psi+\nu)_{,2,2}+\nu^{2}_{,2}+\psi^{2}_{,2}+\psi_{,2}\nu_{,2}$
$-(\psi+\nu)_{,2}\mu_{2,2}\big{)}-\frac{1}{4}e^{2(\psi-\nu-\mu_{2})}q_{,2}^{2}+\frac{1}{4}e^{2(\psi-\nu-\mu_{3})}q_{,3}^{2},$ (35)
$G^{0}{}_{1}=$ $-\frac{1}{2}e^{-(2\psi+\mu_{2}+\mu_{3})}\left((e^{3\psi-\nu+\mu_{3}-\mu_{2}}q_{,2})_{,2}+(e^{3\psi-\nu+\mu_{2}-\mu_{3}}q_{,3})_{,3}\right),$ (36)
$G^{2}{}_{3}=$ $-e^{-\mu_{2}-\mu_{3}}\left((\psi+\nu)_{,2,3}+\nu_{,2}\nu_{,3}+\psi_{,2}\psi_{,3}-(\psi+\nu)_{,2}\mu_{2,3}-(\psi+\nu)_{,3}\mu_{3,2}\right)$
$+\frac{1}{2}e^{2(\psi-\nu)-\mu_{2}-\mu_{3}}q_{,2}q_{,3}.$ (37)
@@
IV The electromagnetic field in stationary axisymmetric space-time
@@
$T^{(m)}{}^{0}{}_{0}$ $=-T^{(m)}{}^{1}{}_{1}=-\frac{1}{8\pi}\bigg{(}e^{-2\psi-2\mu_{2}}(A^{2}_{,2}+B^{2}_{,2})+e^{-2\psi-2\mu_{3}}(A^{2}_{,3}+B^{2}_{,3})\bigg{)},$ (48)
$T^{(m)}{}^{2}{}_{2}$ $=-T^{(m)}{}^{3}{}_{3}=\frac{1}{8\pi}\bigg{(}e^{-2\psi-2\mu_{2}}(A^{2}_{,2}+B^{2}_{,2})-e^{-2\psi-2\mu_{3}}(A^{2}_{,3}+B^{2}_{,3})\bigg{)},$ (49)
@@
V The ansätz of Stückelberg fields $\phi^{a}$
In the original dRGT model [6], the Stückelberg field $\phi^{a}$ are chosen to be unitary gauge $\phi^{a}=x^{\alpha}\delta_{\alpha}^{a}$ and the reference metric is $f_{ab}=\eta_{ab}$. There exists a new basic invariant $I^{ab}=g^{\alpha\beta}\partial_{\alpha}\phi^{a}\partial_{\beta}\phi^{b}$ in massive gravity. When searching the black hole solution, $g^{\mu\nu}$ would possess coordinate singularity in the event horizon. In the unitary gauge, $I^{ab}$ would also possess singularity in the event horizon. The singularities in $I^{ab}$ of a black hole encounter a problem of fluctuations [13]. Thus, in order to obtain the healthy SSS black hole, we need to abandon the requirement of the unitary gauge. In the previous Ref. [14], we use the nonunitary gauge to search for SSS solutions in dRGT theory. When serching the rotating black hole solution, one may also use the nonunitary gauge. However, instead of choosing the nonunitary gauge, we found a more concise approach.

### v3

@@ Abstract
pacs:
04.20.Jbdiscribing text of that key and 04.50.Kddiscribing text of that key and 04.70.Bwdiscribing text of that key and 02.30.Jrdiscribing text of that key
offprints:
1 Introduction
@@
The introduction of a mass term for gravity presents a formidable challenge. The linear mass term is plagued by the van Dam-Veltman-Zakharov (vDVZ) discontinuity VanDam1970 . On the other hand, the nonlinear mass term may propagate the sixth mode on a curved background, known as the Boulware-Deser (BD) ghost Boulware1972 . To circumvent these issues, Comelli and his cooperators try to introduce mass terms without Lorentz symmetry Comelli2012 ; Comelli2013 . We presented a detailed study of the spherically symmetric solutions in Lorentz breaking massive gravity Li2016-2 . Concurrently, de Rham, Gabadadze, and Tolley proposed a healthy Lorentz invariant massive gravity model deRham2011 in 2011.
@@
In dRGT theory, the leading contribution to the decoupling limit is a total derivative which prevents the emergence of ghosts. Beyond the decoupling limit, the lapse function is a Lagrange multiplier that constraint the sixth mode in an arbitrary background. Thus, the dRGT theory described a massive spin-2 particles that propagate 5 healthy degrees of freedom. A comprehensive review of massive gravity is available in deRham2014 . If graviton possesses mass, the gravitational waves would travel at a speed different from light. The recent observation of GW150914 by LIGO Baker2017 has placed an upper bound on the graviton mass $m_{g}\leq 1.2\times 10^{-22}eV$. For more about the graviton mass bounds, you may see Ref. deRham2017 . Fortunately, the dRGT theory can produce a small graviton mass technically natural deRham2013 .
@@
Numerous studies have explored black hole solutions within the dRGT theory. Nieuwenhuizen obtained the Sch-warzschild-de Sitter and Reissner-Nordström-de Sitter black holes Nieuwenhuizen2011 . Cai et. al. also derived the static charged black hole solutions Cai2013 . De Rham and his colleagues Berezhiani2011 highlighted the existence of a new basic invariant $I^{ab}=g^{\alpha\beta}\partial_{\alpha}\phi^{a}\partial_{\beta}\phi^{b}$, whose singularities could pose a problem for fluctuations. In the unitary gauge, it is evident that $I^{ab}=g^{\alpha\beta}\delta_{\alpha}^{a}\delta_{\beta}^{b}$ would exhibit a singularity if $g_{\alpha\beta}$ contains an event horizon. Consequently, there are no self-consistent static spherically symmetric (SSS) solutions in the unitary gauge. In our previous work Li2016-1 , we utilized the nonunitary gauge to search for SSS solutions and discovered seven solutions, including black hole with the Stückelberg hair. We also derive the Vaidya solution in Ref. Li2016-3 and investigated the superradiant instabilities for charged furry black holes in dRGT theory Li2020 . Many papers have also studied the dRGT theory with degenerate reference metrics Cai2015 ; Adams2015 ; Xu2015 ; Cao2016 ; zhang2016 ; zhang2017 . And Ref. Jafari2017 calculate the SSS solutions in arbitrary dimensions.
@@
Recently, we extended Chandrasekhar’s method of calculating rotating black holes into $f(R)$ theory Li2022 ; Li2024-1 ; Li2024-2 . In this paper, we further extend this analytical method into the dRGT massive gravity. We proved that arbitrary nonunitary gauge $\phi^{a}$ with the Minkowski reference metric $\eta_{ab}$ equal to the unitary gauge $\bar{\phi}^{a}=x^{\alpha}\delta^{a}_{\alpha}$ with a curved reference metric $\bar{f}_{ab}$. Thus, instead of chosen the ansätz of the Stückelberg field $\phi^{a}$, we directly choose the axisymmetric reference metric $f_{ab}$ with a unitary gauge $\phi^{a}=x^{\alpha}\delta^{a}_{\alpha}$. The field equations with the electromagnetic fields can be simplified to the Ernst equations. The shortest and simplest route to obtain rotating solutions is to use the conjugate variables. We solve the homogeneous equations and the non-homogeneous equations with the conjugate variables respectively. And the solution that beyond the Kerr-Newman family is obtained.
@@
This paper is organized as follows: Sec. II review the field equation of dRGT theory deRham2011 . In Sec. III, we give a general stationary axisymmetric metric and calculate the Einstein tensor $G_{ab}$. In Sec. IV, we discussed the electromagnetic field equations in the stationary axisymmetric background and the energy-momentum of electromagnetic fields. In Sec. V, we proved that arbitrary nonunitary gauge $\phi^{a}$ with the Minkowski reference metric $\eta_{ab}$ equal to the unitary gauge $\bar{\phi}^{a}=x^{\alpha}\delta^{a}_{\alpha}$ with a curved reference metric $\bar{f}_{ab}$. And we derive the matrix square root $\gamma$ by using the Cayley-Hamilton theorem and calculate the $T^{(\mathcal{K})}_{ab}$. In Sec. VI, we simplify the metric to a highly symmetric form by choosing the gauge freedom. In Sec. VII, we derive the Ernst equations. In Sec. VIII, we transform the variables into a conjugate metric. In Sec. IX, we solve the homogeneous equations of the Ernst equations. In Sec. X, we obtain the non-homogeneous solutions. The last section is a detailed conclusion and discussion.
@@
2 The modified Einstein equations in dRGT theory
@@
And the trace of $\gamma^{i}{}^{\alpha}{}_{\beta}$ is defined as $\gamma^{i}]=g^{\alpha\beta}\gamma^{i}_{\alpha\beta}$. In the original Ref. deRham2011 , the matrix $\gamma^{2}$ is defined by the expression
@@
and $T^{(m)}_{\alpha\beta}$ represents the energy-momentum tensor of mater. It should be pointed out that the following calculations are acually using the tetrad components $a$, not $\alpha$. Due to the Bianchi identity $\nabla_{\alpha}G^{\alpha\beta}=0$ and the conservation of energy-momentum tensor $\nabla_{\alpha}T^{(m)}{}^{\alpha\beta}=0$, there is $\nabla_{\alpha}T^{(\mathcal{K})}{}^{\alpha\beta}=0$. As is pointed in Ref.Cao2016 , the variation of $\gamma^{i}]$ is given by the relationship
@@
[…]
$-2\alpha_{4}((6-6[\gamma]+3\mathcal{U}_{2}-\mathcal{U}_{3})\gamma^{\alpha}{}_{\beta}+3(2-2[\gamma]+\mathcal{U}_{2})\gamma^{2}{}^{\alpha}{}_{\beta}$
[…]
@@
3 Stationary axisymmetric spacetime
@@
$\pm e^{2\mu}[(dx^{2})^{2}+(dx^{3})^{2}],$ (28)
which is shown in Ref. Chandrasekhar:1983 . At last, we shall write the stationary axisymmetric metric in the form
$ds^{2}=$ $-e^{2\nu(x^{2},x^{3})}dt^{2}+e^{2\psi(x^{2},x^{3})}(d\varphi-q(x^{2},x^{3})dt)^{2}$
$+e^{2\mu_{2}(x^{2},x^{3})}(dx^{2})^{2}+e^{2\mu_{3}(x^{2},x^{3})}(dx^{3})^{2},$ (29)
@@
where $\mathrm{d}$ is the exterior differential operator and $\wedge$ stands for Cartan wedge. $\Omega^{a}{}_{b}$ are curvature 2-forms. After direct and tedious mathematical calculation Li2022 ; Li2024-1 ; Li2024-2 , one obtain nonzero components of $G_{ab}$ as follows:
$G^{0}{}_{0}$= $e^{-2\mu_{2}}\left(\psi_{,2,2}+\psi^{2}_{,2}+\psi_{,2}(\mu_{3}-\mu_{2})_{,2}\right)+e^{-2\mu_{3}}\big{(}\psi_{,3,3}$
$+\psi^{2}_{,3}+\psi_{,3}(\mu_{2}-\mu_{3})_{,3}\big{)}+e^{-\mu_{2}-\mu_{3}}\big{(}(e^{\mu_{2}-\mu_{3}}\mu_{2,3})_{,3}$
$+(e^{\mu_{3}-\mu_{2}}\mu_{3,2})_{,2}\big{)}+\frac{1}{4}e^{2(\psi-\nu-\mu_{2})}q_{,2}^{2}+\frac{1}{4}e^{2(\psi-\nu-\mu_{3})}q_{,3}^{2},$ (32)
$G^{1}{}_{1}$= $e^{-2\mu_{2}}\left(\nu_{,2,2}+\nu^{2}_{,2}+\nu_{,2}(\mu_{3}-\mu_{2})_{,2}\right)+e^{-2\mu_{3}}\big{(}\nu_{,3,3}$
$+\nu^{2}_{,3}+\nu_{,3}(\mu_{2}-\mu_{3})_{,3}\big{)}+e^{-\mu_{2}-\mu_{3}}\big{(}(e^{\mu_{2}-\mu_{3}}\mu_{2,3})_{,3}$
$+(e^{\mu_{3}-\mu_{2}}\mu_{3,2})_{,2}\big{)}-\frac{3}{4}e^{2(\psi-\nu-\mu_{2})}q_{,2}^{2}-\frac{3}{4}e^{2(\psi-\nu-\mu_{3})}q_{,3}^{2},$ (33)
$G^{2}{}_{2}$= $e^{-2\mu_{2}}\left(\psi_{,2}\nu_{,2}+(\psi+\nu)_{,2}\mu_{3,2}\right)+e^{-2\mu_{3}}\big{(}(\psi+\nu)_{,3,3}$
$+\nu^{2}_{,3}+\psi^{2}_{,3}+\psi_{,3}\nu_{,3}-(\psi+\nu)_{,3}\mu_{3,3}\big{)}$
$+\frac{1}{4}e^{2(\psi-\nu-\mu_{2})}q_{,2}^{2}-\frac{1}{4}e^{2(\psi-\nu-\mu_{3})}q_{,3}^{2},$ (34)
$G^{3}{}_{3}$= $e^{-2\mu_{3}}\left(\psi_{,3}\nu_{,3}+(\psi+\nu)_{,3}\mu_{2,3}\right)+e^{-2\mu_{2}}\big{(}(\psi+\nu)_{,2,2}$
$+\nu^{2}_{,2}+\psi^{2}_{,2}+\psi_{,2}\nu_{,2}-(\psi+\nu)_{,2}\mu_{2,2}\big{)}-\frac{1}{4}e^{2(\psi-\nu-\mu_{2})}q_{,2}^{2}$
$+\frac{1}{4}e^{2(\psi-\nu-\mu_{3})}q_{,3}^{2},$ (35)
$G^{0}{}_{1}$= $-\frac{1}{2}e^{-(2\psi+\mu_{2}+\mu_{3})}\big{(}(e^{3\psi-\nu+\mu_{3}-\mu_{2}}q_{,2})_{,2}$
$+(e^{3\psi-\nu+\mu_{2}-\mu_{3}}q_{,3})_{,3}\big{)},$ (36)
$G^{2}{}_{3}$= $-e^{-\mu_{2}-\mu_{3}}\big{(}(\psi+\nu)_{,2,3}+\nu_{,2}\nu_{,3}+\psi_{,2}\psi_{,3}-(\psi+\nu)_{,2}\mu_{2,3}$
$-(\psi+\nu)_{,3}\mu_{3,2}\big{)}+\frac{1}{2}e^{2(\psi-\nu)-\mu_{2}-\mu_{3}}q_{,2}q_{,3}.$ (37)
@@
4 The electromagnetic field in stationary axisymmetric space-time
@@
$T^{(m)}{}^{0}{}_{0}$ $=-T^{(m)}{}^{1}{}_{1}=-\frac{1}{8\pi}\bigg{(}e^{-2\psi-2\mu_{2}}(A^{2}_{,2}+B^{2}_{,2})$
$+e^{-2\psi-2\mu_{3}}(A^{2}_{,3}+B^{2}_{,3})\bigg{)},$ (48)
$T^{(m)}{}^{2}{}_{2}$ $=-T^{(m)}{}^{3}{}_{3}=\frac{1}{8\pi}\bigg{(}e^{-2\psi-2\mu_{2}}(A^{2}_{,2}+B^{2}_{,2})$
$-e^{-2\psi-2\mu_{3}}(A^{2}_{,3}+B^{2}_{,3})\bigg{)},$ (49)
@@
5 The ansätz of Stückelberg fields $\phi^{a}$
In the original dRGT model deRham2011 , the Stückelberg field $\phi^{a}$ are chosen to be unitary gauge $\phi^{a}=x^{\alpha}\delta_{\alpha}^{a}$ and the reference metric is $f_{ab}=\eta_{ab}$. There exists a new basic invariant $I^{ab}=g^{\alpha\beta}\partial_{\alpha}\phi^{a}\partial_{\beta}\phi^{b}$ in massive gravity. When searching the black hole solution, $g^{\mu\nu}$ would possess coordinate singularity in the event horizon. In the unitary gauge, $I^{ab}$ would also possess singularity in the event horizon. The singularities in $I^{ab}$ of a black hole encounter a problem of fluctuations Berezhiani2011 . Thus, in order to obtain the healthy SSS black hole, we need to abandon the requirement of the unitary gauge. In the previous Ref. Li2016-1 , we use the nonunitary gauge to search for SSS solutions in dRGT theory. When serching the rotating black hole solution, one may also use the nonunitary gauge. However, instead of choosing the nonunitary gauge, we found a more concise approach.